OpenStudy (anonymous):

A cage at a pet store has 9 white mice. In how many different ways can Dana choose 3 white mice to buy?

OpenStudy (anonymous):

(n!/(n-r)!r!)= (9!/(9-3)!3!) (9!/(6)!3!) = cancel 6 from top and bottom (9!8!7!)/3!2! = 504/6 84 ways